12. The equation y = 10x – 70 describes the amount of money a class of students might earn from candy bar sales. What are the slope and y-intercept of this line?
Answer by josmiceli(19441)

When the equation has the form y = mx + b, the slope is m and the y intercept is b.
slope = 10
y intercept = -70
Note that the y intercept occurs when x = 0.
Make x = 0 in the equation to check.
You can verify that m is really the slope, too.
y = 10x - 70
Call a particular x and y that solve the eqation (x[1], y[1])
Call another x and y that solve it (x[2], y[2])
So now you can write
y[2] = 10x[2] - 70
y[1] = 10x[1] - 70
Now subtract the second equation from the first
y[2] - y[1] = 10(x[2] - x[1])
y[2] - y[1] / x[2] - x[1] = 10
This is the definition of slope, and it does equal 10 or the m in y = mx + b.